# A Basic Evaluation of Neural Networks Trained with the Error Diffusion Learning Algorithm

### v2

@@ 4.1 Parity Check Task
To evaluate the efficacy of the Error Diffusion Learning Algorithm (EDLA), we initially assessed its performance on a parity check task, a classical benchmark in neural network learning. This task involves determining whether the number of ones in a binary input vector is even or odd. The parity check task is a two-class classification problem. The EDLA network consists of input, hidden, and output layers. The neurons in the hidden layers use the sigmoid or ReLU activation functions (AFs). The output neuron is singular and employs the sigmoid AF, producing a value between 0 and 1. The output is interpreted as the probability of the input vector containing an odd number of ones. Thus, the output neuron is expected to yield a value above 0.5 for odd counts and below 0.5 for even counts. The total number of possible input combinations is $2^{n_{\mathrm{bit}}}$, where $n_{\mathrm{bit}}$ denotes the number of bits in the input vector. In this experiment, $n_{\mathrm{bit}}=2,3,4,5$ with each dataset including all possible combinations for a given $n_{\text{bit}}$. A mini-batch size of 4 is consistently employed across all configurations.
@@
Figure 2 illustrates the performance of the EDLA network with a single hidden layer and the sigmoid AF across various learning rates (0.1, 1.0, 10.0) and neuron counts per layer (8, 16, 32, 64, 128). The results demonstrate that the EDLA network achieves high accuracy on the parity check task. The network attains perfect accuracy across all cases except for configurations with eight neurons when $n_{\mathrm{bit}}=5$. The number of epochs required to reach an accuracy of 0.9 decreases as the number of neurons increases. For learning rates of 10 and 1.0, the convergence time (reach epoch) is smaller than the learning rate of 0.1. Furthermore, for $n_{\mathrm{bit}}\in\{3,5\}$, the reach epoch at a learning rate of 1.0 is nearly identical to that at a learning rate of 10. For $n_{\mathrm{bit}}=2$, the reach-epoch for a learning rate of 1.0 is larger than that at a learning rate of 10, but for $n_{\mathrm{bit}}=4$ the reach epoch at a learning rate of 1.0 is smaller than that at a learning rate of 10. These results suggest that learning rates of 10 and 1.0 enable efficient training.
@@
Figure 3 presents an analysis of the performance of the EDLA network using the sigmoid AF across various hidden layers (1, 2, 4, and 8 hidden layers) and neuron counts per layer with a fixed learning rate of 1.0. The results reveal that the accuracy consistently approaches or reaches 1.0 across different bit numbers ($n_{\mathrm{bit}}$), regardless of the number of hidden layers. However, the number of epochs required to achieve 0.9 accuracy reveals more nuanced dependencies on the network depth and neuron count. Specifically, for $n_{\mathrm{bit}}=2$, shallow networks (one layer) exhibit a relatively low reach epoch for smaller neuron numbers, but the reach epoch increased for larger neuron counts. The deeper networks do not exhibit significant reach epoch reductions for $n_{\mathrm{bit}}=2$. Conversely, for $n_{\mathrm{bit}}\in\{3,4\}$ with deeper architectures (four and eight layers), the reach epoch increases with the number of neurons. For $n_{\mathrm{bit}}=5$, the reach epoch decreases notably with the neuron count and stabilized for neuron counts of 64 and above. These results indicate that EDLA networks with sigmoid AF can consistently achieve high accuracy on the parity check task, with optimal performance depending on the complexity of the task and the chosen network configuration.
@@
Figure 5 shows the accuracy and convergence speed of the EDLA network with the ReLU across various neuron counts and hidden layer depths. Both the accuracy and convergence speed improve with increasing neuron counts in the ReLU network. However, this improvement in the convergence speed does not continue indefinitely and eventually plateaus. These results indicate a performance comparable to that of sigmoid-based networks, with a consistently faster convergence observed for ReLU-based configurations. This enhanced convergence speed arises from the absence of an upper bound in the ReLU, facilitating larger weight updates.
@@
Figure 2: Performance of EDLA with the sigmoid AF across various learning rates on the parity check task. The left column depicts the accuracy of the EDLA network with the sigmoid AF as a function of neurons in a single hidden layer for different learning rates (LR = 0.1, 1.0, 10.0). The accuracies are averaged from 19900 to 20000 epochs. The right column shows the number of epochs required to reach an accuracy of 0.9 as a function of neurons in a hidden layer for the same learning rates. Each row of plots corresponds to a different bit length (n_bit, ranging from 2 to 5), for the parity check task. The results are averaged over ten independent trials.
@@
Figure 3: Accuracy and convergence characteristics of the EDLA network with the sigmoid AF across various network depths (1, 2, 4, and 8 hidden layers) at a fixed learning rate of 1.0. The left column shows the average classification accuracy as a function of the neurons per hidden layer. The accuracies are average from 19900 to 20000 epochs. The right column presents the number of epochs required to reach an accuracy of 0.9 as a function of neurons per hidden layer. Each row of plots corresponds to different bit lengths (n_bit, ranging from 2 to 5). The results are averaged over ten independent trials.
@@
Figure 4: Performance of EDLA with the ReLU AF across various learning rates in parity check task. The left column shows the accuracy of the EDLA network as a function of neurons in a single hidden layer for different learning rates (LR = 0.01, 0.1, 1.0). The accuracies are average from 19900 to 20000 epochs. The right column shows the number of epochs required to achieve an accuracy of 0.9 against the number of neurons in the hidden layer for the same learning rates. Each row of plots corresponds to a different bit number (n_bit), ranging from 2 to 5, for the parity check task. Missing data points in the left column indicate that the accuracy is excessively low. Missing data points in the right column indicate that the accuracy does not reach 0.9. The results are averaged over ten independent trials.
@@
Figure 5: Accuracy and convergence characteristics of the EDLA network with ReLU AF across various network depths (1, 2, 4, and 8 hidden layers), using a fixed learning rate of 1.0. The left column shows average classification accuracy as a function of neurons per hidden layer. The accuracies are average from 19900 to 20000 epochs. The right column presents the number of epochs required to reach an accuracy of 0.9 as a function of the number of neurons per hidden layer. Each row of plots corresponds to different bit lengths (n_bit), ranging from 2 to 5. The results are averaged over ten independent trials.
@@ 4.2 Regression Tasks
The EDLA network consists of an input, hidden, and output layer. The neurons in the hidden layers use the sigmoid or ReLU AFs. The neurons in the output layer use the linear AF (i.e., no AF). The training parameters are carefully selected for better performance, with a learning rate of 0.001 for the sigmoid AF and 0.0001 for the ReLU AF. Training is conducted for 500 epochs employing a mini-batch size of 64 samples.
Figure 6 illustrates the Mean Absolute Error (MAE) performance of the EDLA network across varying numbers of neurons (8, 16, 32, 64, 128, 256, 512, and 1024) and different layer configurations (1, 2, and 4 layers). The results indicate that EDLA networks using sigmoid AF with one and two layers achieve efficient performance across almost all neuron-count configurations for all datasets. The EDLA network, which has a single hidden layer using the ReLU AF, also achieves efficient performance. This improves performance as the number of neurons in the hidden layers increases. Additionally, increasing the network depth does not yield performance enhancements and exhibits deterioration compared to shallower network architectures. This deterioration will be caused by the instability of the EDLA network due to the increase in both neuron count and network depth. These observations collectively suggest that, while the EDLA approach is adequately capable of regression tasks, augmenting the number of neurons and network depth does not necessarily improve performance.
Figure 6: Performance of the EDLA network on regression datasets across varying layer configurations. The left column shows the Mean Absolute Error (MAE) for sigmoid AF, while the right column displays the MAE for the ReLU AF. Both sets of results are plotted as a function of the number of neurons per hidden layer for 1, 2, and 4 hidden layers. Each row corresponds to the Airfoil Self-Noise dataset (top row), Concrete Compressive Strength dataset (middle row), and Energy Efficiency dataset (bottom row). Missing data points in the figure signify that the network output becomes excessively large or yields a not-a-number (NaN) value. The results are averaged over four independent trials.
@@ 4.3 Image Classification Tasks
To evaluate the performance of the EDLA network in image classification tasks, the experiments are conducted using the handwritten digits dataset (Digits), MNIST, and Fashion MNIST datasets. The Digits used consists of 5620 grayscale images with a resolution of $8\times 8$ pixels. Because the Digits dataset is not divided into training and test datasets, 20 % of the entire dataset is randomly selected and used as test data. The MNIST comprises 60,000 training and 10,000 testing grayscale images of handwritten digits, each with a resolution of $28\times 28$ pixels. Similarly, the Fashion-MNIST includes 60,000 training and 10,000 testing grayscale images of fashion articles at a resolution of $28\times 28$ pixels.
@@
The EDLA network consists of an input layer, a hidden layer, and an output layer. The neurons in the input and hidden layers utilize the sigmoid or ReLU AFs. The output layer employs the sigmoid AF. To assess its performance across various configurations, the number of neurons in each hidden layer is varied systematically. The learning rates when using the sigmoid and ReLU AFs are 1.0 and 0.1, respectively. The network is trained for 500 epochs. We employed the minibatch method and set the minibatch size to 128 samples.
@@
Figure 7 illustrates the performance of the EDLA network using the sigmoid AF for various numbers of neurons in each hidden layer. Across all datasets, the classification performance generally improved as the number of neurons in the hidden layers increased. Specifically, for the Digits and MNIST datasets, near-perfect accuracy is rapidly achieved, even with a relatively small number of neurons. For the Fashion MNIST datasets, while high accuracy is attained, a larger number of neurons contributed to an incrementally improved performance. However, the depth of the EDLA network does not improve the performance, but worsens the performance. In terms of the reach epoch, for all configurations, the EDLA network shows a sufficiently fast reach of 0.8 accuracy. While deeper networks will lead to a slower reach, the number of neurons in each hidden layer will lead to a faster reach.
@@
Similarly, Figure 8 shows the performance of the EDLA network using the ReLU AF across varying numbers of neurons in each hidden layer. The results reveal that the EDLA network with the ReLU AF achieves a performance comparable to that of a sigmoid-based network. However, for the ReLU AF, deeper networks lead to worse performance. Furthermore, for a deeper network, the number of neurons in each hidden layer leads to worse performance. The reach epoch for the ReLU AF decreases with the number of neurons in each hidden layer. However, the reach epoch for deeper networks will lead to slower reach.
@@
Figure 7: Performance of EDLA with the sigmoid AF on image datasets across layer configurations. The plots in the left column show the accuracy of the EDLA network with the sigmoid AF as a function of the number of neurons in each hidden layer for different layer configurations (1, 2, and 4 layers) across three datasets: digits, MNIST, and fashion_MNIST. The plots in the right column illustrate the number of epochs required to achieve an accuracy of 0.8 against the number of neurons in each hidden layer for the same layer configurations. The results are averaged over four independent trials.
@@
Figure 8: Performance of EDLA with the ReLU AF on image datasets across layer configurations. The plots in the left column show the accuracy of the EDLA network with the sigmoid AF as a function of the number of neurons in each hidden layer for different layer configurations (1, 2, and 4 layers) across three datasets: digits, MNIST, and fashion_MNIST. The plots in the right column illustrate the number of epochs required to achieve an accuracy of 0.8 against the number of neurons in each hidden layer for the same layer configurations. The results are averaged over four independent trials.

### v3

@@ 4.1 Parity Check Task
To evaluate the efficacy of the Error Diffusion Learning Algorithm (EDLA), we initially assessed its performance on a parity check task, a classic benchmark in neural network learning. This task involves determining whether the number of ones in a binary input vector is even or odd. The parity check task is a two-class classification problem. The EDLA network consists of input, hidden, and output layers. The neurons in the hidden layers use the sigmoid or ReLU activation functions (AFs). The network has a single-output neuron that employs the sigmoid AF, producing a value between 0 and 1. The output is interpreted as the probability of the input vector containing an odd number of ones. Thus, the output neuron is expected to yield a value above 0.5 for odd counts and below 0.5 for even counts. The total number of possible input combinations is $2^{n_{\mathrm{bit}}}$, where $n_{\mathrm{bit}}$ denotes the number of bits in the input vector. In this experiment, $n_{\mathrm{bit}}=2,3,4,5$ with each dataset including all possible combinations for a given $n_{\text{bit}}$. A mini-batch size of 4 is consistently employed across all configurations.
@@
Figure 2 illustrates the performance of the EDLA network with a single hidden layer and the sigmoid AF across various learning rates (0.1, 1.0, 10.0) and neuron counts per layer (8, 16, 32, 64, 128). The results demonstrate that the EDLA network achieves high accuracy on the parity check task. The network attains perfect accuracy across all cases except for configurations with eight neurons when $n_{\mathrm{bit}}=5$. The epochs required to reach an accuracy of 0.9 decrease as the number of neurons increase. For learning rates of 10 and 1.0, the convergence time (reach epoch) is shorter than that for a learning rate of 0.1. Furthermore, for $n_{\mathrm{bit}}\in\{3,5\}$, the reach epoch at a learning rate of 1.0 is nearly identical to that at a learning rate of 10. For $n_{\mathrm{bit}}=2$, the reach epoch for a learning rate of 1.0 is larger than that at a learning rate of 10, but for $n_{\mathrm{bit}}=4$ the reach epoch at a learning rate of 1.0 is smaller than that at a learning rate of 10. These results suggest that learning rates of 10 and 1.0 enable efficient training.
@@
Figure 3 presents an analysis of the performance of the EDLA network using the sigmoid AF across various hidden layers (1, 2, 4, and 8 hidden layers) and neuron counts per layer with a fixed learning rate of 1.0. The results reveal that the accuracy consistently approaches or reaches 1.0 across different bit numbers ($n_{\mathrm{bit}}$), regardless of the number of hidden layers. However, the number of epochs required to achieve 0.9 accuracy reveals more nuanced dependencies on the network depth and neuron count. Specifically, for $n_{\mathrm{bit}}=2$, shallow networks (one layer) exhibit a relatively low reach epoch for smaller neuron numbers, but the reach epoch increases for larger neuron counts. The deeper networks do not exhibit significant reach epoch reductions for $n_{\mathrm{bit}}=2$. Conversely, for $n_{\mathrm{bit}}\in\{3,4\}$ with deeper architectures (four and eight layers), the reach epoch increases with the number of neurons. For $n_{\mathrm{bit}}=5$, the reach epoch decreases notably with the neuron count and stabilizes for neuron counts of 64 and above. These results indicate that EDLA networks with the sigmoid AF can consistently achieve high accuracy on the parity check task, with optimal performance depending on the complexity of the task and the chosen network configuration.
@@
Figure 5 shows the accuracy and convergence speed of the EDLA network with the ReLU across various neuron counts and hidden layer depths for a learning rate of 0.1. Both the accuracy and convergence speed improve with increasing neuron counts in the ReLU network. However, this improvement in the convergence speed does not continue indefinitely and eventually plateaus. These results indicate a performance comparable to that of sigmoid-based networks, with a consistently faster convergence observed for ReLU-based configurations. This enhanced convergence speed arises from the absence of an upper bound in the ReLU, facilitating larger weight updates.
@@
Figure 2: Performance of EDLA with the sigmoid AF across various learning rates on the parity check task. The left column depicts the accuracy of the EDLA network with the sigmoid AF as a function of neurons in a single hidden layer for different learning rates (LR = 0.1, 1.0, 10.0). The accuracies are averaged from 19900 to 20000 epochs. The right column shows the number of epochs required to reach an accuracy of 0.9 as a function of the number of neurons in a hidden layer for the same learning rates. Each row of plots corresponds to a different bit length (n_bit, ranging from 2 to 5), for the parity check task. The results are averaged over ten independent trials.
@@
Figure 3: Accuracy and convergence characteristics of the EDLA network with the sigmoid AF across various network depths (1, 2, 4, and 8 hidden layers) at a fixed learning rate of 1.0. The left column shows the average classification accuracy as a function of the neurons per hidden layer. The accuracies are averaged from 19900 to 20000 epochs. The right column presents the number of epochs required to reach an accuracy of 0.9 as a function of the number of neurons per hidden layer. Each row of plots corresponds to different bit lengths (n_bit, ranging from 2 to 5). The results are averaged over ten independent trials.
@@
Figure 4: Performance of EDLA with the ReLU AF across various learning rates in the parity check task. The left column shows the accuracy of the EDLA network as a function of neurons in a single hidden layer for different learning rates (LR = 0.01, 0.1, 1.0). The accuracies are averaged from 19900 to 20000 epochs. The right column shows the number of epochs required to achieve an accuracy of 0.9 against the number of neurons in the hidden layer for the same learning rates. Each row of plots corresponds to a different bit length (n_bit), ranging from 2 to 5, for the parity check task. Missing data points in the left column indicate that the accuracy is excessively low. Missing data points in the right column indicate that the accuracy does not reach 0.9. The results are averaged over ten independent trials.
@@
Figure 5: Accuracy and convergence characteristics of the EDLA network with ReLU AF across various network depths (1, 2, 4, and 8 hidden layers), using a fixed learning rate of 0.1. The left column shows average classification accuracy as a function of neurons per hidden layer. The accuracies are averaged from 19900 to 20000 epochs. The right column presents the number of epochs required to reach an accuracy of 0.9 as a function of the number of neurons per hidden layer. Each row of plots corresponds to different bit lengths (n_bit), ranging from 2 to 5. The results are averaged over ten independent trials.
@@ 4.2 Regression Tasks
The EDLA network consists of an input layer, hidden layers, and an output layer. The neurons in the hidden layers use the sigmoid or ReLU AFs. The neurons in the output layer use the identity AF (i.e., no AF). The training parameters are carefully selected for better performance, with a learning rate of 0.001 for the sigmoid AF and 0.0001 for the ReLU AF. Training is conducted for 500 epochs employing a mini-batch size of 64 samples.
Figure 6 illustrates the Mean Absolute Error (MAE) performance of the EDLA network across varying numbers of neurons (8, 16, 32, 64, 128, 256, 512, and 1024) and different layer configurations (1, 2, and 4 layers). The results indicate that EDLA networks using sigmoid AF with one and four layers achieve low MAE across almost all neuron-count configurations for all datasets. However, the EDLA network with the sigmoid AF does not exhibit performance improvements as the number of neurons or network depth increases.
Notably, the EDLA network with a single hidden layer using the ReLU AF not only achieves low MAE but also consistently outperforms the 1-layer sigmoid network across all three datasets. Furthermore, for this 1-layer network, performance slightly improves as the number of neurons increases. Additionally, increasing the network depth does not yield performance improvements but rather leads to performance degradation compared with shallower architectures. This deterioration is likely caused by the instability of the EDLA network when using ReLU, due to the increase in both neuron count and network depth. As detailed in the caption (Figure 6), this instability was particularly severe for the 4-layer ReLU network, resulting in catastrophically large MAE values (e.g., $10^{30}$) or NaN values for several configurations
These observations collectively suggest that, while the EDLA approach is adequately capable of handling regression tasks, augmenting the number of neurons and network depth does not necessarily improve performance. Furthermore, the EDLA network with the ReLU AF tends to exhibit inferior performance with increased depth. Despite this instability in deeper networks, the ReLU AF demonstrated a slight performance advantage over the sigmoid AF in the most stable (1-layer) configurations.
Figure 6: Performance of the EDLA network on regression tasks across varying layer configurations. Mean Absolute Error (MAE) is shown against the number of neurons per hidden layer for networks with 1, 2, and 4 hidden layers. The rows correspond to the Airfoil Self-Noise (top), Concrete Compressive Strength (middle), and Energy Efficiency (bottom) datasets. The left column shows results for networks using the sigmoid AF (learning rate = 0.001), while the right column shows results for ReLU activation (learning rate = 0.0001). Each point represents the average of four independent trials. Missing data points for the 4-layer ReLU network indicate training instability, where the MAE became excessively large or a NaN (not-a-number) value. The omitted data points (number of neurons, MAE) are as follows: Airfoil: (64, $4.39\times 10^{3}$), (128, $2.66\times 10^{14}$), (256, $1.90\times 10^{31}$), (512, NaN), (1024, NaN); Concrete: (64, $3.09\times 10^{3}$), (128, $3.13\times 10^{13}$), (256, $9.84\times 10^{31}$), (512, $5.79\times 10^{19}$), (1024, NaN); Energy: (64, $1.84\times 10^{4}$), (128, $1.07\times 10^{20}$), (256, NaN), (512, NaN), (1024, NaN).
@@ 4.3 Image Classification Tasks
To evaluate the performance of the EDLA network in image classification tasks, the experiments are conducted using the Digits dataset (Optical Recognition of Handwritten Digits), MNIST, Fashion-MNIST, and CIFAR-10 datasets. The Digits dataset consists of 5620 grayscale images with a resolution of $8\times 8$ pixels. Because the Digits dataset is not divided into training and test datasets, 20 % of the entire dataset is randomly selected and used as test data. The MNIST comprises 60,000 training and 10,000 testing grayscale images of handwritten digits, each with a resolution of $28\times 28$ pixels. Similarly, the Fashion-MNIST includes 60,000 training and 10,000 testing grayscale images of Fashion articles at a resolution of $28\times 28$ pixels. The CIFAR-10 consists of 50,000 training and 10,000 testing color images with a resolution of $32\times 32$ pixels. The CIFAR-10 dataset is more complex than the other three datasets because it contains color images. In this experiment, the CIFAR-10 dataset is converted to a flattened one-dimensional vector format.
@@
The EDLA network consists of an input layer, hidden layers, and an output layer. The neurons in the hidden layers use either the sigmoid or ReLU activation functions. The output layer employs the sigmoid AF. To assess its performance across various configurations, the number of neurons in each hidden layer is varied systematically. The learning rates when using the sigmoid and ReLU AFs are 1.0 and 0.1, respectively. The network is trained for 500 epochs when using Digits, MNIST, and Fashion-MNIST datasets, and for 1000 epochs when using the CIFAR-10 dataset. A mini-batch size of 128 samples is used.
@@
Figure 7 illustrates the performance of the EDLA network using the sigmoid AF for various numbers of neurons in each hidden layer. Across all datasets, the classification performance generally improved as the number of neurons in the hidden layers increased. Specifically, for the Digits and MNIST datasets, near-perfect accuracy is rapidly achieved, even with a relatively small number of neurons. For the Fashion-MNIST dataset, while high accuracy is attained, a larger number of neurons contributed to incrementally improved performance. For the CIFAR-10 dataset, the performance is lower compared to the other datasets. This lower performance of CIFAR-10 is attributed to the network architecture, which has a simple layer structure and is not well-suited for processing complex color images as in convolutional neural networks. The performance improvement with an increasing number of neurons is more pronounced in the CIFAR-10 dataset. However, the depth of the EDLA network does not improve performance and can even degrade it. In terms of the reach epoch, for all configurations, the EDLA network shows a sufficiently fast reach of 0.2 and 0.8 accuracy when using the CIFAR-10 dataset and the others. While deeper networks will lead to a slower reach, the number of neurons in each hidden layer will lead to a faster reach.
@@
Similarly, Figure 8 shows the performance of the EDLA network using the ReLU AF across varying numbers of neurons in each hidden layer. The results reveal that the EDLA network with the ReLU AF achieves a performance comparable to that of the sigmoid-based network. However, similar to the sigmoid network, increasing network depth generally leads to worse performance. The trend of accuracy versus neuron count is also more complex: while performance on CIFAR-10 generally improves with more neurons, performance on Digits (4-layer) and Fashion-MNIST (2-layer) slightly degrades. The reach epoch for the ReLU AF also decreases with the number of neurons in each hidden layer. The reach epoch for the ReLU AF is larger (slower) than that for the sigmoid AF.
These observations collectively suggest that EDLA networks can effectively handle relatively simple image classification tasks using both sigmoid and ReLU activation functions. However, more complex datasets such as CIFAR-10 present challenges for the simple EDLA architecture employed in this study. Increasing the number of neurons in each hidden layer generally enhances performance, while increasing network depth does not necessarily yield improvements and may even degrade performance. Regarding convergence speed, increasing the number of neurons generally accelerated convergence for both activation functions. Conversely, increasing network depth tended to slow convergence for sigmoid networks. Furthermore, the reach epoch for the ReLU AF was generally larger (slower) than that for the sigmoid AF across most configurations.
@@
Figure 7: Performance of EDLA with the sigmoid AF on image datasets across layer configurations. The plots in the left column show the accuracy of the EDLA network with the sigmoid AF as a function of the number of neurons in each hidden layer for different layer configurations (1, 2, and 4 layers) across four datasets: Digits, MNIST, Fashion-MNIST, and CIFAR-10. The plots in the right column illustrate the number of epochs required to achieve an accuracy of $\theta_{\mathrm{threshold}}$ against the number of neurons in each hidden layer for the same layer configurations. For CIFAR-10, $\theta_{\mathrm{threshold}}$ is set to 0.2. Otherwise, $\theta_{\mathrm{threshold}}$ is set to 0.8. The results are averaged over four independent trials.
@@
Figure 8: Performance of EDLA with the ReLU AF on image datasets across layer configurations. The plots in the left column show the accuracy of the EDLA network with the ReLU AF as a function of the number of neurons in each hidden layer for different layer configurations (1, 2, and 4 layers) across four datasets: Digits, MNIST, Fashion-MNIST, and CIFAR-10. The plots in the right column illustrate the number of epochs required to achieve an accuracy of $\theta_{\mathrm{threshold}}$ against the number of neurons in each hidden layer for the same layer configurations. For CIFAR-10, $\theta_{\mathrm{threshold}}$ is set to 0.2. Otherwise, $\theta_{\mathrm{threshold}}$ is set to 0.8. The results are averaged over four independent trials.